OpenStudy (anonymous):

The principal P is borrowed at a simple interest rate r for a period of time t. Find the loan's future value A, or the total amount due at time t. Round answer to the nearest cent. P = $6,000, r = 6%, t = 2 years

OpenStudy (campbell_st):

The simple interest formula is \[I = P \times \frac{r}{100} \times n\] this will give the amount of interest owed. the total to be repaid, of future value A = Principal (P) + Interest(I) hope it helps

OpenStudy (anonymous):

that doesn't help to much haha can you go through the steps with me?

OpenStudy (campbell_st):

ok... the interest on the loan is \[I = 6000 \times \frac{6}{100} \times 2\] calculate that

OpenStudy (anonymous):

so 720?

OpenStudy (campbell_st):

ok... so thats the interest you have to pay on the loan. you have to repay the amount borrowed ($6000) plus the interest ($720) so the future value A = 6000 + 720

OpenStudy (anonymous):

can you help me with this one too? Calculate the monthly payments for an add-on interest loan of $3,500, with an annual interest rate of 19% and a term of 1.5 years. Round to the nearest cent.

OpenStudy (campbell_st):

Will it use the simple interest formula..?

OpenStudy (anonymous):

I'm not sure.. :(